OpenStudy (anonymous):

Find the Vertex: f(x)= -3x^2-7x+6

OpenStudy (anonymous):

@saifoo.khan helpppp lol

OpenStudy (saifoo.khan):

Hey. To find vertex there's a shortcut.

OpenStudy (anonymous):

uh huh

OpenStudy (saifoo.khan):

to find the x-cordinate, use this formula: -b/2a

OpenStudy (saifoo.khan):

-(-7)/2(-3)

OpenStudy (anonymous):

that will give me the vertex ?

OpenStudy (saifoo.khan):

Yes. "x-cordinate" of vertex.

OpenStudy (saifoo.khan):

Solve it first..

OpenStudy (anonymous):

kk hold on :)

OpenStudy (saifoo.khan):

Holding.. :)

OpenStudy (anonymous):

got 7/6

OpenStudy (anonymous):

vertex is (h,k) right? so i believe that is the "h" ?

OpenStudy (saifoo.khan):

Perfect. ^

OpenStudy (saifoo.khan):

Find k now.

OpenStudy (anonymous):

how to find k lol ?

OpenStudy (saifoo.khan):

Lol. Use GOogle. Jk.

OpenStudy (anonymous):

i got it lmao :))

OpenStudy (anonymous):

hold on

OpenStudy (saifoo.khan):

Okay. so now use 7/6 to find it. y = -3x^2-7x+6 y = -3(7/6)^2 -7(7/6) + 6 solve.

OpenStudy (anonymous):

got decimal :( hate decimal

OpenStudy (anonymous):

so the vertex would be : (7/6, -6.24) ?

OpenStudy (saifoo.khan):

Should be 121/12

OpenStudy (anonymous):

wtf ?

OpenStudy (anonymous):

how? :O

OpenStudy (anonymous):

-156/25

OpenStudy (anonymous):

hey come back here >:o

OpenStudy (saifoo.khan):

VERTEX SHOULD BE -7/6. -_-

OpenStudy (saifoo.khan):

THE X-CORDINATE.

OpenStudy (anonymous):

CALM DOWN !! CAPS WAR XD

OpenStudy (saifoo.khan):

ONE NEGATIVE SIGN RUINS EVERYTHING. >_<

OpenStudy (anonymous):

so -7/6 because its opposite of 7/6 ?

OpenStudy (anonymous):

you suppose to change it everytime ?

OpenStudy (anonymous):

how come my teacher didn't change it -_____-

OpenStudy (saifoo.khan):

No. we calculated the x-cordinate wrong in the first place.

OpenStudy (anonymous):

oh....XD

OpenStudy (saifoo.khan):

\[-\frac{b}{2a} \to -\frac{(-7)}{2(-3)}\to \frac{7}{-6} \to -\frac 76\]

OpenStudy (anonymous):

ohhhhhhhhh i see now lmao hhahaha :D

OpenStudy (anonymous):

omfg...my whole problem is wrong, thanks saif ! lol : XD

OpenStudy (anonymous):

+99 medals

OpenStudy (saifoo.khan):

;D Thanks.

OpenStudy (anonymous):

hmm instead of 121/10 should i just leave it as 10.08 cuz i have to graph this afterward ?

OpenStudy (anonymous):

great now i gotta calculate x-int and y-int :| let see......

OpenStudy (saifoo.khan):

Sure. Same thing. Depends on in what format your teacher wants

OpenStudy (anonymous):

stay there.. lol

OpenStudy (saifoo.khan):

Lol. okays.

OpenStudy (anonymous):

stuck lol....find x-intercept?

OpenStudy (anonymous):

i know let y=0

OpenStudy (anonymous):

0=-3x^2-7x+6

OpenStudy (anonymous):

-6=-3x^2-7x <-----stuck here..

OpenStudy (saifoo.khan):

Use quadratic formula

OpenStudy (anonymous):

oh yee

OpenStudy (anonymous):

but the way my teacher did it was to set y=0 and just add and divide stuff on left and right side...which is confusing but i suppose we can use quadratic aswell :)

OpenStudy (saifoo.khan):

Your teacher must have used the "completing the square" method. It's hella confusing. It confuses me too. :/

OpenStudy (anonymous):

yeah thats what that was, completing the square lol.... ok i got y= 7+- sqrt(-23)/-6 and what's next ?

OpenStudy (anonymous):

i dont think u can reduce it down or anything right? and u can't even square root a negative

OpenStudy (saifoo.khan):

No. We can't reduce.

OpenStudy (saifoo.khan):

u can't even square root a negative Yes you're very right.

OpenStudy (anonymous):

ohh i get it, it's 7+- 23i/-6

OpenStudy (saifoo.khan):

Roots are: -14 and 0.

OpenStudy (saifoo.khan):

Roots are: -3 and 2/3

OpenStudy (anonymous):

what? @_@

OpenStudy (anonymous):

was that you on the video? https://www.khanacademy.org/math/algebra/quadratics/quadratic_formula/v/quadratic-formula-1

OpenStudy (saifoo.khan):

Haha. Ofcourse not.